An object has a mass of 300 g. (a) What is its weight on Earth? (b) What is its mass on the Moon? (c) What will be its acceleration on the Moon when a 0.500 N resultant force acts on it?


1
Expert's answer
2020-07-20T14:56:49-0400

(a) The weighi of the object is:


"W = mg"

where "m" is its mass and "g = 9.81 m\/s^2" is the gravitational acceleration. Thus:


"W = 0.3kg\cdot 9.81 m/s^2 = 2.943 N"

(b) The mass of any object on the Moon is the same as on Earth. In this case 300 g.

(c) The acceleration can be calculated using the Newton's second law:


"a = \dfrac{F}{m}"

where "F = 0.5N" is the resultant force. Thus, obtain:


"a = \dfrac{0.5N}{0.3 kg} \approx 1.67 m/s^2"

Answer. a) 2.943 N, b) 300 g, c)1.67 m/s^2.